OpenStudy (anonymous):

On the graph of the equation 5x – 2y = 20, what is the value of the y-intercept?

OpenStudy (anonymous):

Put x = 0 y = -10

OpenStudy (anonymous):

At y-intercept, x is always 0. So plug in 0 into your equation for x. 5(0) - 2y = 20 -2y = 20 y = -10